OpenStudy (anonymous):

Write an equation in slope-intercept form for each line described here: containing (0,-4) with the slope of 8

OpenStudy (anonymous):

y-(-4)=8(x-0), y=8x-4

OpenStudy (anonymous):

how did you get that?